# Probing the Decision Boundaries of In-context Learning in Large Language Models

### v2

@@ Abstract
In-context learning is a key paradigm in large language models (LLMs) that enables them to generalize to new tasks and domains by simply prompting these models with a few exemplars without explicit parameter updates. Many attempts have been made to understand in-context learning in LLMs as a function of model scale, pretraining data, and other factors. In this work, we propose a new mechanism to probe and understand in-context learning from the lens of decision boundaries for in-context binary classification. Decision boundaries are straightforward to visualize and provide important information about the qualitative behavior of the inductive biases of standard classifiers. To our surprise, we find that the decision boundaries learned by current LLMs in simple binary classification tasks are often irregular and non-smooth, regardless of linear separability in the underlying task. This paper investigates the factors influencing these decision boundaries and explores methods to enhance their generalizability. We assess various approaches, including training-free and fine-tuning methods for LLMs, the impact of model architecture, and the effectiveness of active prompting techniques for smoothing decision boundaries in a data-efficient manner. Our findings provide a deeper understanding of in-context learning dynamics and offer practical improvements for enhancing robustness and generalizability of in-context learning. [^1] Preprint. Our code is released at https://github.com/siyan-zhao/ICL_decision_boundary.
@@ 4 Experiments
Classification Tasks. We investigate the decision boundary of LLMs by prompting them with $n$ in-context examples of binary classification tasks, with an equal number of examples for each class. We generate classification datasets using scikit-learn (Pedregosa et al., 2011), creating three types of linear and non-linear classification tasks: linear, circle, and moon, each describing different shapes of ground-truth decision boundaries. Detailed information on the dataset generation can be found in Appendix H. In addition to the in-context examples, we calculate the in-context learning accuracy on a held-out test set of size 100. We sample in-context examples and test points from classification task and convert them into prompt, with an example shown in Appendix G.
@@
Figure 2 compares the decision boundaries of 6 LLMs when provided with 128 in-context examples. Even on simple linearly separable classification problems, all of these models exhibit non-smooth decision boundaries. The decision boundaries vary significantly across models, indicating that these models have different reasoning abilities to interpret the same in-context data. All models show fragmented decision regions, which means small changes in the input features can result in different classifications. This raises concerns about the reliability of LLMs and their practical deployment, as even when test accuracy for classification is high (shown in Figure 3, where test accuracy increases with the number of context examples), the underlying decision boundary lacks generalization. In the following sections, we will explore factors that affect decision boundary smoothness and investigate methods to improve smoothness.
@@ 4.3 How to Improve the Decision Boundary Smoothness?
Can We Finetune LLMs on the In-Context Examples to Achieve Smoother Decision Boundaries? Our experiments indicate that finetuning LLMs on in-context examples does not result in smoother decision boundaries. Specifically, we finetuned Llama3-8B on 128 in-context learning examples and found that the resulting decision boundaries remained non-smooth. Examples of the decision boundaries after finetuning are provided in Appendix C.
@@
where the expectation is with respect to task, data points $\{(x_{i},y_{i})\}_{i=1}^{N}$, and the number of context points $m$. After training, we evaluate the same finetuned model on various binary classification tasks with varying numbers of context points. To ensure the test tasks are unseen during training, we use different parameters in creating the datasets, such as the separateness between two classes and the scale between the inner and outer circles in the circle task. See Appendix H for more details.
@@
How to Use Uncertainty-aware Active Learning to Learn Decision Boundaries We investigate whether the decision boundary can be smoothed by providing the LLM with labels of the most uncertain points on the grid as additional in-context examples. Uncertainty is measured as the entropy of the probability distribution of the two classes after softmax normalization of the logits. Our study focuses on an active learning scheme where new in-context examples are incrementally added based on the LLM’s current uncertainty. Initially, we obtain the decision boundary conditioned on the existing in-context examples. To refine this boundary, we query the LLM over a grid and select the top-k most uncertain points, ensuring they are spatially distant from each other using a greedy sampling approach. For labeling these uncertain points, we use a logistic regression model well-trained on a larger dataset with perfect accuracy as the ground truth decision boundary. As shown in Figure 11, this uncertainty-aware active sampling method results in a smoother decision boundary over iterations compared to random sampling. The iterative refinement enhances the model’s generalization capabilities, leading to higher test set accuracies and greater sample efficiency, requiring fewer additional in-context examples to achieve performance gains. These findings indicate that leveraging the LLM’s uncertainty measurements is valuable for selecting new in-context examples in resource-constrained settings where labeled data is scarce. We show more examples in Appendix F.
@@ 5 Related Works
Learning to learn in-context In contrast to the emergent in-context capabilities of LLMs, existing works have also studied methods that learn to perform in-context learning explicitly. Min et al. (2022a) propose MetaICL, a meta-training framework for finetuning pretrained LLMs to perform in-context learning on a large and diverse collection of tasks. MetaICL outperforms several baselines including emergent in-context learning and multi-task learning followed by zero-shot transfer. Going beyond the text domain, TNP (Nguyen and Grover, 2022) and PFNs (Müller et al., 2021) are two concurrent works that propose to train transformer models to perform in-context prediction for a family of functions, which allows in-context generalization to unseen functions after training. Similarly, (Garg et al., 2022) show that autoregressive transformers can be trained from scratch to learn function classes such as linear functions and 2-layer ReLU networks. Zhao et al. (2023) also propose meta-learning transformers to in-context learn group preferences, serving as an in-context learned reward model that adapts to diverse group preferences. These works present an interesting set of baselines for our work to examine the in-context learning ability of LLMs.
@@ Appendix A Pretrained LLMs decision boundary on linear and non-linear classification tasks
Appendix B SFT LLMs for in-context classification
@@ Appendix E Traditional Classifiers Model Details
Appendix F Uncertainty Aware Active Sampling For Smoother Decision Boundary and Better Test set Accuracy
(a) Active sampling
(b) Random sampling
Figure 16: Comparison of active and random sampling methods.
(a) Active sampling
(b) Random sampling
Figure 17: Comparison of active and random sampling methods.
Appendix G Prompt Format for binary classification
@@
Appendix H Classification Datasets Creation Details
@@
Appendix I Limitation
One limitation of our study is the focus on demonstrating mainly binary classification tasks. Limited by the available compute, we chose binary tasks and also for better qualitative reasoning. However, we also extended our experiments to tasks with four classes and found that our methods generalize to multi-class classification and other more complex tasks. Additionally, the exploration of fine-tuning and adaptive sampling methods, although effective in our experiments, may not be universally applicable across closed-source LLMs that do not allow access to logits. Future work should consider a broader range of tasks and datasets, as well as a more diverse set of LLM models, to validate and extend our findings.

### v3

@@ Abstract
In-context learning is a key paradigm in large language models (LLMs) that enables them to generalize to new tasks and domains by simply prompting these models with a few exemplars without explicit parameter updates. Many attempts have been made to understand in-context learning in LLMs as a function of model scale, pretraining data, and other factors. In this work, we propose a new mechanism to probe and understand in-context learning from the lens of decision boundaries for in-context binary classification. Decision boundaries are straightforward to visualize and provide important information about the qualitative behavior of the inductive biases of standard classifiers. To our surprise, we find that the decision boundaries learned by current LLMs in simple binary classification tasks are often irregular and non-smooth, regardless of linear separability in the underlying task. This paper investigates the factors influencing these decision boundaries and explores methods to enhance their generalizability. We assess various approaches, including training-free and fine-tuning methods for LLMs, the impact of model architecture, and the effectiveness of active prompting techniques for smoothing decision boundaries in a data-efficient manner. Our findings provide a deeper understanding of in-context learning dynamics and offer practical improvements for enhancing robustness and generalizability of in-context learning. [^1] Our code is released at https://github.com/siyan-zhao/ICL_decision_boundary.
@@ 4 Experiments
Classification Tasks. We investigate the decision boundary of LLMs by prompting them with $n$ in-context examples of binary classification tasks, with an equal number of examples for each class. We generate classification datasets using scikit-learn (Pedregosa et al., 2011), creating three types of linear and non-linear classification tasks: linear, circle, and moon, each describing different shapes of ground-truth decision boundaries. Detailed information on the dataset generation can be found in Appendix G. In addition to the in-context examples, we calculate the in-context learning accuracy on a held-out test set of size 100. We sample in-context examples and test points from classification task and convert them into prompt, with an example shown in Appendix F.
@@
Figure 2 compares the decision boundaries of 6 LLMs when provided with 128 in-context examples. Even on simple linearly separable classification problems, all of these models exhibit non-smooth decision boundaries. The decision boundaries vary significantly across models, indicating that these models have different reasoning abilities to interpret the same in-context data. All models show fragmented decision regions, which means small changes in the input features can result in different classifications. This raises concerns about the reliability of LLMs and their practical deployment, as even when test accuracy for classification is high (shown in Figure 3, where test accuracy increases with the number of context examples), the underlying decision boundary lacks generalization. We further demonstrate nonsmoothness in NLP text classification tasks by projecting text input into 2D space, as detailed in Appendix H. In the following sections, we will explore factors that affect decision boundary smoothness and investigate methods to improve smoothness.
@@ 4.3 How to Improve the Decision Boundary Smoothness?
Can We Finetune LLMs on the In-Context Examples to Achieve Smoother Decision Boundaries? Our experiments indicate that finetuning LLMs on in-context examples does not result in smoother decision boundaries. Specifically, we finetuned Llama3-8B on 128 in-context learning examples and found that the resulting decision boundaries remained non-smooth. Examples of the decision boundaries after finetuning are provided in Appendix B.
@@
where the expectation is with respect to task, data points $\{(x_{i},y_{i})\}_{i=1}^{N}$, and the number of context points $m$. After training, we evaluate the same finetuned model on various binary classification tasks with varying numbers of context points. To ensure the test tasks are unseen during training, we use different parameters in creating the datasets, such as the separateness between two classes and the scale between the inner and outer circles in the circle task. See Appendix G for more details.
@@
How to Use Uncertainty-aware Active Learning to Learn Decision Boundaries We investigate whether the decision boundary can be smoothed by providing the LLM with labels of the most uncertain points on the grid as additional in-context examples. Uncertainty is measured as the entropy of the probability distribution of the two classes after softmax normalization of the logits. Our study focuses on an active learning scheme where new in-context examples are incrementally added based on the LLM’s current uncertainty. Initially, we obtain the decision boundary conditioned on the existing in-context examples. To refine this boundary, we query the LLM over a grid and select the top-k most uncertain points, ensuring they are spatially distant from each other using a greedy sampling approach. For labeling these uncertain points, we use a logistic regression model well-trained on a larger dataset with perfect accuracy as the ground truth decision boundary. As shown in Figure 11, this uncertainty-aware active sampling method results in a smoother decision boundary over iterations compared to random sampling. The iterative refinement enhances the model’s generalization capabilities, leading to higher test set accuracies and greater sample efficiency, requiring fewer additional in-context examples to achieve performance gains. These findings indicate that leveraging the LLM’s uncertainty measurements is valuable for selecting new in-context examples in resource-constrained settings where labeled data is scarce. We show more examples in Appendix I.
@@ 5 Related Works
Learning to learn in-context In contrast to the emergent in-context capabilities of LLMs, existing works have also studied methods that learn to perform in-context learning explicitly. Min et al. (2022a) propose MetaICL, a meta-training framework for finetuning pretrained LLMs to perform in-context learning on a large and diverse collection of tasks. MetaICL outperforms several baselines including emergent in-context learning and multi-task learning followed by zero-shot transfer. Beyond text, TNP (Nguyen and Grover, 2022; Nguyen et al., 2023; Nguyen and Grover, 2024) and PFNs (Müller et al., 2021) propose to train transformer models to perform in-context prediction for a family of functions, which allows in-context generalization to unseen functions after training. Similarly, Garg et al. (2022) show that autoregressive transformers can be trained from scratch to learn function classes such as linear functions and 2-layer ReLU networks. Other work also shows that alignment can be done in-context (Zhao et al., 2023), where in-context learned reward model can be used for inference-time preference alignment. These works present an interesting set of baselines for our work to examine the in-context learning ability of LLMs.
@@ 6 Conclusion
Acknowledgments
This research is supported by NSF CAREER Award 2341040, Schmidt Sciences AI2050 Fellowship, Samsung, and Cisco.
@@ Appendix A Pretrained LLMs decision boundary on linear and non-linear classification tasks
Appendix B Finetune on in-context examples only
Appendix C SFT LLMs for in-context classification
@@ Appendix E Traditional Classifiers Model Details
Appendix F Prompt Format for binary classification
@@
Appendix G Classification Datasets
@@
Appendix H Decision Boundary of LLMs on NLP tasks.
We extend our analysis to multi-class NLP classification tasks using high-dimensional real-world datasets. To address the challenge of visualizing high-dimensional text embeddings, we project them onto a 2D space using t-SNE and send the 2D embeddings as input in the prompt to the LLM. While any dimensionality reduction technique inevitably introduces confounding factors, this approach allows us to extend our analysis to more complex, real-world scenarios. Our experiments encompass six widely-used NLP classification tasks, covering both binary and multi-class settings. These include Subjective/Obejective sentence classification (SUBJ) (Conneau and Kiela, 2018), financial sentiment analysis (FP) (Malo et al., 2014), textual entailment recognition (RTE) (Wang et al., 2019), hate speech detection (ETHOS) (Mollas et al., 2020), sentiment analysis (SST-2) (Socher et al., 2013) and news topic classification (AG_NEWS) (Zhang et al., 2015). The results, presented in Figure 17, demonstrate that the non-smooth decision boundary characteristics observed in our synthetic datasets persist in these more complex NLP tasks.
Figure 17: Decision boundaries of Llama-3-8b on six NLP tasks, ranging from binary to multi-class classification. Since text embeddings are natively high-dimensional, we projected text embeddings onto a 2D space using t-SNE. The irregular, non-smooth behaviors are also seen in these tasks.
Appendix I Uncertainty Aware Active Sampling For Smoother Decision Boundary and Better Test set Accuracy
(a) Active sampling
(b) Random sampling
Figure 18: Comparison of active and random sampling methods.
(a) Active sampling
(b) Random sampling
Figure 19: Comparison of active and random sampling methods.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: It reflects our proposed method and experimental findings.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide limitation in the appendix section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical studies.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We present complete method and experiment details in Section 3 and 4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All code and checkpoints will be released publicly upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We present complete method and experiment details in Section 3 and 4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Although we do not provide error bar for every plot, this is due to the nature of our work, since we are visualizing the decision boundary for qualitative understanding. We justify this with additional plots in various settings in the appendix. Apart from the decision boundary plots, we do plot the accuracy plot with error bar.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the compute in the appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work does not have societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: NA
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes] ,
  59. Justification: Yes, we cited every dataset and models we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We don’t release new dataset.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: we don’t have any.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We don’t have this.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.